OpenStudy (anonymous):

How many solutions are there to the equation sec(100x) + 4 =-0.2 satisfying 0<= x <= 2pi

OpenStudy (anonymous):

sec 100x = -4.2 cos100 x = 1/sec100x =1/-4.2 = -0.2381 100x = 180-76.23 = 103.77 degrees (least value)